OpenStudy (anonymous):

If \(\Large \frac{a}{b+c} +\frac{b}{a+c}+\frac{c}{a+b}=1\), then fine \(\Large \frac{a^2}{b+c} +\frac{b^2}{a+c}+\frac{c^2}{a+b}\)

OpenStudy (anonymous):

Gotcha! :D

OpenStudy (anonymous):

Let \(\large 1=\frac{a+b+c}{a+b+c} \) Take \(\large a+b+c\) of the denominotor to the LHS and multiply with every term: \(\Large \frac{a(a+b+c)}{b+c} +\frac{b(a+b+c)}{a+c} + \frac{c(a+b+c)}{a+b}=a+b+c\) Then split the terms according to the requirements of the question and you will get: \(\Large \frac{a^2}{b+c}+\frac{b^2}{a+c}+\frac{c^2}{a+b}+\frac{a(b+c)}{b+c}+\frac{b(a+c)}{a+c}+\frac{c(a+b)}{a+b}=a+b+c\\\large \Rightarrow \frac{a^2}{b+c}+\frac{b^2}{a+c}+\frac{c^2}{a+b}=a+b+c-(a+b+c)=0\)

OpenStudy (kainui):

How did you come up with the idea to let 1=(a+b+c)/(a+b+c) to be useful to you?

OpenStudy (kainui):

At any rate, this is awesome, thanks haha.